# Effective Heterogeneous Federated Learning via Efficient Hypernetwork-based Weight Generation

### v1

@@ 5.1. Experiment Setup
Dataset and Model. In this work, we use three different datasets for our evaluations and a VGG-based baseline CNN model architecture (Simonyan and Zisserman, 2014). We discuss the details of the dataset and model used as follows.
SVHN dataset (Netzer et al., 2011) includes 99,289 labeled images of 10 digits (1-10) extracted from real-world house numbers, offering a diverse representation of physical world images and is well-suited for mobile image classification tasks. The model comprises four convolutional blocks, each with two layers featuring channels (32, 64, 128, 256), followed by max-pooling at the end of each block.
STL10 dataset (Coates et al., 2011) comprises 10 classes with diverse real-world objects, containing 1.3K samples per class. Images in STL10 are high-resolution at 96x96 pixels, distinguishing it from datasets like CIFAR, making STL10 suitable for real-world evaluations of HypeMeFed in federated learning. The model architecture mirrors that used for SVHN, but with channel configurations of 64, 128, 256, and 512.
UniMiB SHAR dataset (Micucci et al., 2017) comprises 4.2K samples from 30 subjects, capturing data from smartphone accelerometers for eight fall types. UniMiB SHAR represents a widely adopted real-world sensing application. We adopted the same model configuration as the STL10 dataset.
@@
HypeMeFed integrates three exit layers into the federated learning model. To accommodate heterogeneous clients, we categorize 50 clients into three groups: 17 using a small model with one exit layer, another 17 using a medium-capacity model with two exit layers, and the remaining 16 using the full model with all three exit layers. For UniMiB, we configure 30 clients (10 per group) due to its limited dataset size. To assure heterogeneity in the training process, 20% of the clients were uniformly selected from the three groups to participate in each federated learning round. Except for FedAvg, which does not support device heterogeneity, all other baselines maintained model capacities balanced across client groups. For LotteryFL, we relaxed resource constraints initially to fully train the model for its anticipated performance.
@@ 5.2. Overall Accuracy Performance
Next, while HypeMeFed targets to address device heterogeneity, we examine its robustness against data heterogeneity by analyzing the accuracy with varying non-IID degrees in the data (). Figure 8 depicts the accuracy for four different values: 0.1 (strong non-IID), 0.5, 1.0, and 10,000 (IID). The results suggest that as data becomes uniformly distributed, accuracy tends to improve. However, most baselines still fail to outperform FedAvg-S, while HypeMeFed consistently achieves better or at least competitive model accuracy.
@@ 5.3.2. Weight Generation for Batch Norm and Fully Connected Layers
We hypothesize that this counter-intuitive observation arises from the increased number of denominators in the aggregation process due to the generated weights, thereby reducing the impact of model personalization-related information (e.g., locally trained BN and FC layers). In contrast, the feature extractor (convolutional layers) focuses on extracting universally applicable features across client models (Lubana et al., 2022; Xu et al., 2021; Hu et al., 2023), maximizing the benefits of weight generation.
@@ 5.3.3. Impact of Rank on SVD Operations
To validate this observation, Figure 12 shows the Mean Absolute Error (MAE) of the weights for different . Quantitatively, the results confirm that as increases, the hypernetwork-generated weights closely approximate the original, with the scenario where the hypernetwork generates all weights showing the lowest MAE. This indicates that leveraging a hypernetwork with an appropriate effectively balances computational efficiency and model accuracy, thereby supporting robust federated learning across heterogeneous clients.
@@ 5.4. Applicability to Large Models
Experiments conducted so far utilize a relatively small-sized model with a limited number of convolutional layers. The next question we answer is whether HypeMeFed can scale to larger and deeper models. For this, we examine the performance with (i) a more complex model with extensive parameters and (ii) models with deeper layers. Specifically, to test a model with extensive parameters, we conducted experiments with ResNet18 (He et al., 2016)—a relatively large model with 11.5M parameters, featuring a residual connection architecture. We used the CIFAR100 dataset (Krizhevsky et al., 2009), containing 50,000 image samples across 100 categories. Three intermediate multi-exit layers were added to the baseline ResNet18 model, splitting it depth-wise into three segments.
Figure 13 plots the global model accuracy trend with increasing federated learning rounds. The results suggest that HypeMeFed is applicable to larger models, as it successfully converges and shows improved model accuracy compared to the baseline where all users utilize the smallest model (FedAvg-S in Fig. 13). In fact, HypeMeFed achieves performance comparable to the upper bound assuming all clients can utilize the largest model (i.e., FedAvg-L). Here, training the full-rank hypernetwork for ResNet18 requires more than 44.74 GB (5.59B parameters) of memory, which is infeasible and computationally extensive. However, with =100 for the hypernetwork optimizations, only 113.28 MB (14M parameters) was needed, reducing the memory usage by 99.87%. These results demonstrate that HypeMeFed can effectively scale to larger models while maintaining high accuracy and significantly reducing memory overhead.
@@
Figure 14 (a) plots the normalized Mean Absolute Error (MAE) between the original and generated weights with respect to layer depth. Note that we present normalized MAE given that the magnitude of the weight norm differs across layers. As the results show, the normalized MAE tends to increase with layer depth despite generating all weights (=Full). This trend likely arises from the challenge of accurately capturing global features in deeper layers, as discussed in previous studies (Luo et al., 2021; Lee et al., 2024b). Despite this observation, HypeMeFed achieves a higher global accuracy of 89.47% (+3.71%) compared to the FedAvg-S (85.76%), where all clients use the lightest (1-exit) model. Figure 14 (b) plots the MAE between original and generated weights at the final convolution layer. Given an auto-regressive generation approach, where we predict layer weights using layer weights, we assess the accumulated error by comparing MAE across different initial layers used to generate the final layer. Notice that the MAE remains consistent across initiating layers, indicating the robustness of the proposed hypernetwork.
@@ 5.5. Evaluation on Real-World Testbed
To demonstrate the practicality of HypeMeFed in real-world scenarios, we evaluated HypeMeFed on a testbed of heterogeneous embedded platforms: 4 Raspberry Pi 4, 4 Nvidia Jetson Nano, and 4 NVIDIA Jetson TX2, totaling 12 devices with varying computational resources—weak, moderate, and ample, respectively. Additionally, we used a server with an Nvidia RTX 3090 GPU, an Intel i9-K@3.6GHz CPU, and 64GB RAM. We test the UniMiB dataset with a VGG model, where each client holds a non-overlapping dataset distributed following the Dirichlet distribution (=0.5). FedAvg-S and FedAvg-L universally used single- and three-layered models, respectively, while clients in HypeMeFed used models suitable for each device’s resources. Unless otherwise specified, all configurations matched those mentioned in Section 5.
Figure 15 plots the latency per federated learning round for different client platforms along with the server-side latency (left) and the global model accuracy for FedAvg-L, FedAvg-S, and HypeMeFed (right). With all clients using the full model, FedAvg-L achieves high accuracy, but the Raspberry Pi shows prolonged latency in processing this model, potentially delaying the overall federated learning process. In FedAvg-S, latency is kept low, but using a single-layer model causes a 20.41% accuracy drop. By applying HypeMeFed and adaptively configuring models for each device, we balance training latency and accuracy, outperforming FedAvg-S by 5.13%. Note a naive hypernetwork induces 3.9 sec of server computation time and 5.6 GB of memory overhead, while our optimizations (=100) reduce this to 2.1 sec (1.86 speedup) and 100.8 MB (98.22% reduction), agreeing with our previous results. Overall, our testbed validations confirm that HypeMeFed can be a practical and efficient framework for supporting federated learning on heterogeneous clients.
@@ 6. Discussions
Optimizing hypernetwork architectures and performance. In this work, we demonstrated the feasibility of exploiting hypernetworks in federated learning through HypeMeFed. While HypeMeFed effectively addresses the complexity issues by leveraging SVD-based low-rank factorization, research on hypernetworks is still in its early stages. We identify two potential research directions. First, HypeMeFed utilizes the same hypernetwork architecture across all layers. Exploring optimal configurations for each layer may further enhance hypernetwork performance. Second, despite reduced complexity, training hypernetworks introduces server-side latency, delaying the federated learning process. Developing efficient hypernetwork training algorithms or suitable alternatives could mitigate this issue and improve overall system performance.
Diversifying model architectures. In this work, we focused our evaluations on CNN-based architectures. However, our experiences show that HypeMeFed can also support other architectures, such as MLPs, RNNs, and LSTMs, by effectively generating weights using its hypernetwork. As we expand our research to include more complex neural network architectures, such as Transformers, we anticipate that the complexity of the hypernetwork will present significant challenges that needs to be addressed.
@@ Acknowledgements.
We thank Jaeho Jin for setting up the testbed and Sinyoung Oh for designing the figures.References

### v2

@@ 5.1. Experiment Setup
Dataset and Model. In this work, we use three different datasets for our evaluations and a VGG-based baseline CNN model architecture (Simonyan and Zisserman, 2014) with small variations to suit each dataset’s characteristics. We discuss the details of the dataset and model used as follows.
SVHN dataset (Netzer et al., 2011) includes 99,289 labeled 3232 pixel images of 10 digits (1-10) extracted from real-world house numbers, offering a diverse representation of physical world images and is well-suited for mobile image classification tasks. The model comprises four convolutional blocks, each with two layers featuring channels (32, 64, 128, 256), followed by max-pooling at the end of each block.
STL10 dataset (Coates et al., 2011) comprises 10 classes with diverse real-world objects, containing 1.3K samples per class. Images in STL10 have a relatively higher resolution of 96x96 pixels compared to the SVHN, allowing us to validate HypeMeFed under different input data conditions. The model architecture mirrors that used for SVHN, but with channel configurations of 64, 128, 256, and 512.
UniMiB SHAR dataset (Micucci et al., 2017) comprises 4.2K samples from 30 subjects, capturing three channel (x, y, z-axis) time-series sensor data from smartphone accelerometers for eight fall types. UniMiB SHAR represents a widely adopted real-world sensing application exploiting time-series sensor data. We adopted the same model configuration as the STL10 dataset.
Note that we choose these datasets to ensure diversity across different real-world scenarios, including image classification tasks with varying levels of complexity, and time-series data for common mobile and embedded sensing applications. The use of such datasets allow us to comprehensively evaluate HypeMeFed in heterogeneous federated learning environments for various potential use cases. Furthermore, while we present results using model architectures based on CNNs, we note that the proposed approach is applicable to different network architectures (c.f., Section 6).
@@
Throughout the evaluation, by default, HypeMeFed integrates three exit layers into the federated learning model. To accommodate heterogeneous clients, we categorize 50 clients into three groups: 17 using a small model with one exit layer, another 17 using a medium-capacity model with two exit layers, and the remaining 16 using the full model with all three exit layers. For UniMiB, we configure 30 clients (10 per group) due to its limited dataset size. We also test HypeMeFed with varying exit counts (more than three) and locations to further examine the feasibility and generalizability of HypeMeFed. To assure heterogeneity in the training process, 20% of the clients were uniformly selected from the three groups to participate in each federated learning round. Except for FedAvg, which does not support device heterogeneity, all other baselines maintained model capacities balanced across client groups. For LotteryFL, we relaxed resource constraints initially to fully train the model for its anticipated performance.
@@ 5.2. Overall Accuracy Performance
Next, while HypeMeFed targets to address device heterogeneity, we examine its robustness against data heterogeneity by analyzing the accuracy with varying non-IID degrees in the data (). Figure 8 depicts the accuracy for four different values: 0.1 (strong non-IID), 0.5, 1.0, and 10,000 (IID). The results suggest that as data becomes uniformly distributed, accuracy tends to improve. However, most baselines still fail to outperform FedAvg-S, while HypeMeFed consistently achieves better or at least competitive model accuracy. We note that the HypeMeFed can coexist with advanced aggregation or local training mechanisms to address the issue of device and/or data heterogeneity.
@@ 5.3.2. Weight Generation for Batch Norm and Fully Connected Layers
We hypothesize that this counter-intuitive observation arises from the increased number of denominators in the aggregation process due to the hypernetwork-generated weights, thereby reducing the impact of model personalization-related information (e.g., locally trained BN and FC layers). In contrast, the feature extractor (i.e., convolutional layers) focuses on extracting universally applicable features across client models (Lubana et al., 2022; Xu et al., 2021; Hu et al., 2023), maximizing the benefits of weight generation.
@@ 5.3.3. Impact of Rank on SVD Operations
To validate this observation, Figure 12 shows the Mean Absolute Error (MAE) of the weights for different . Quantitatively, the results here confirm that as increases, the hypernetwork-generated weights closely approximate their respective original values, with the scenario where the hypernetwork generates all weights showing the lowest MAE. This indicates that leveraging a hypernetwork with an appropriate effectively balances computational efficiency and model accuracy, thereby supporting robust federated learning across heterogeneous clients.
@@ 5.4. Applicability to Large Models
Experiments conducted so far utilize a relatively small-sized model with a limited number of convolutional layers. The next question we answer is whether HypeMeFed can scale to larger and deeper models. Additionally, we also conduct experiments with varying the numbers of exits to explore whether it remains effective when models are split in different granularities. For this, we examine the performance with (i) a more complex model with extensive parameters and (ii) models with deeper layers. Specifically, to test a model with extensive parameters, we conducted experiments with ResNet18 (He et al., 2016)—a relatively large model with 11.5M parameters, featuring a residual connection architecture. We used the CIFAR100 dataset (Krizhevsky et al., 2009), containing 50,000 image samples across 100 categories. Three intermediate multi-exit layers were added to the baseline ResNet18 model, splitting it depth-wise into three segments.
Figure 13 plots the global model accuracy trend with increasing federated learning rounds. The results suggest that HypeMeFed is applicable to larger models, as it successfully converges and shows improved model accuracy compared to the baseline where all users utilize the smallest model (FedAvg-S in Fig. 13). In fact, HypeMeFed achieves performance comparable to the upper bound assuming all clients can utilize the largest model (i.e., FedAvg-L). Here, training the full-rank hypernetwork for ResNet18 requires more than 44.74 GB (5.59B parameters) of memory, which is infeasible and computationally extensive. However, with =100 for the hypernetwork optimizations, only 113.28 MB (14M parameters) was needed, reducing the memory usage by 99.87%. These results demonstrate that HypeMeFed can effectively scale to larger models while maintaining high accuracy and significantly reducing memory overhead, showcasing its applicability in real-world deployments.
@@
Figure 14 (a) plots the normalized Mean Absolute Error (MAE) between the original and generated weights with respect to layer depth. Note that we present normalized MAE given that the magnitude of the weight norm differs across layers.
As the results show, the normalized MAE tends to increase with layer depth despite generating all weights (=Full). This trend likely arises from the challenge of accurately capturing global features in deeper layers, as discussed in previous studies (Luo et al., 2021; Lee et al., 2024b). Despite this observation, HypeMeFed achieves a higher global accuracy of 89.47% (+3.71%) compared to the FedAvg-S (85.76%), where all clients use the lightest (1-exit) model. Figure 14 (b) plots the MAE between original and generated weights at the final convolution layer. Given an auto-regressive generation approach, where we predict layer weights using layer weights, we assess the accumulated error by comparing MAE across different initial layers used to generate the final layer. Notice that the MAE remains consistent across initiating layers, indicating the robustness of the proposed hypernetwork. These results confirm that HypeMeFed’s hypernetwork-based approach effectively handles deeper models and larger networks with more exit points, maintaining both accuracy and computational efficiency.
@@ 5.5. Evaluation on Real-World Testbed
To demonstrate the practicality of HypeMeFed in real-world scenarios, we evaluated HypeMeFed on a testbed of heterogeneous embedded platforms: 4 Raspberry Pi 4, 4 Nvidia Jetson Nano, and 4 NVIDIA Jetson TX2, totaling 12 devices with varying computational resources—weak, moderate, and ample, respectively (c.f. Figure 16). Additionally, we used a server with an Nvidia RTX 3090 GPU, an Intel i9-K@3.6GHz CPU, and 64GB RAM. We test the UniMiB dataset with a VGG model, where each client holds a non-overlapping dataset distributed following the Dirichlet distribution (=0.5). FedAvg-S and FedAvg-L universally used single- and three-layered models, respectively, while clients in HypeMeFed used models suitable for each device’s resources. Unless otherwise specified, all configurations matched those mentioned in Section 5.
Figure 15 plots the latency per federated learning round for different client platforms along with the server-side latency (left) and the global model accuracy for FedAvg-L, FedAvg-S, and HypeMeFed (right). As expected, FedAvg-L achieves high accuracy, but the Raspberry Pi shows prolonged latency in processing this model, potentially delaying the overall federated learning process. In FedAvg-S, latency is kept low, but using a single-layer model causes a 20.41% accuracy drop. By applying HypeMeFed and adaptively configuring models for each device, we balance training latency and accuracy, outperforming FedAvg-S by 5.13%. Note that a naive hypernetwork induces 3.9 sec of server computation time and 5.6 GB of memory overhead, while our optimizations (=100) reduce this to 2.1 sec (1.86 speedup) and 100.8 MB (98.22% reduction), agreeing with our previous results. Overall, our testbed validations confirm that HypeMeFed can be a practical and efficient framework for supporting federated learning on heterogeneous clients.
@@ 6. Discussions
Optimizing hypernetwork architectures and performance. In this work, we demonstrated the feasibility of exploiting hypernetworks in federated learning through HypeMeFed. While HypeMeFed effectively addresses the complexity issues by leveraging SVD-based low-rank factorization, research on hypernetworks is still in its early stages. We identify two potential research directions. First, HypeMeFed utilizes the same hypernetwork architecture across all layers. Exploring optimal configurations for each layer may further enhance hypernetwork performance. Second, despite reduced complexity, training hypernetworks introduces server-side latency, delaying the federated learning process. Developing efficient hypernetwork training algorithms or suitable alternatives could mitigate this issue and improve overall system performance. In addition, to better leverage hypernetworks in device-heterogeneous federated learning and understand their performance, better-defined metrics for evaluating hypernetwork performance may be needed. For example, beyond traditional metrics such as accuracy and MAE (as used in this work), it would be helpful to assess the contextual meaningfulness of the generated weights. Furthermore, evaluation metrics that help identify the sweet spot within the hypernetwork complexity and required resources trade-off could ensure a more seamless integration of hypernetworks into federated learning.
Diversifying model architectures. In this work, we focused our evaluations on CNN-based architectures. However, we emphasize that HypeMeFed is inherently model-agnostic, and its design can be extended to support a wide range of neural network architectures beyond CNNs. Specifically, our hypernetwork-based weight generation approach can be applied to other architectures, such as MLPs, RNNs, and LSTMs, by leveraging the layer-wise relationships within the model to generate missing or incomplete weights. This flexibility allows HypeMeFed to be applicable to various federated learning scenarios involving diverse model types: enabling it to be used in broader applications ranging from natural language processing tasks with RNNs or LSTMs to structured data processing using MLPs. As we expand our research to include more complex neural network architectures, such as Transformers, we anticipate new challenges. For instance, Transformer-based architectures, known for their attention mechanisms and high parameter complexity, may increase the demands on hypernetwork capacity and efficiency. The dynamic and dense inter-dependencies of such models may require more sophisticated hypernetwork designs capable of capturing both local and global relationships across layers, while still minimizing memory and computational overhead. Addressing these challenges could open up new possibilities for deploying HypeMeFed in cutting-edge applications such as NLP, vision-language cooperation applications, and other domains that require sophisticated model architectures. HypeMeFed ’s model-agnostic nature holds significant potential for future federated learning environments, allowing it to adapt to evolving model architectures and application domains.
@@ Acknowledgements.
We thank the anonymous reviewers and shepherd for their efforts in providing constructive comments for the paper. The authors would also like to thank Jaeho Jin for help in configuring the testbed environment and Sinyoung Oh for designing the figures. This work was supported by the Korean Ministry of Science and ICT (MSIT) and IITP under grants IITP-2024-2020-0-01461 (ITRC Program) and IITP-2022-0-00420. Hyung-sin Kim and JeongGil Ko are the corresponding authors for this work.References
@@
- Chauhan et al. (2024) Vinod Kumar Chauhan, Jiandong Zhou, Ping Lu, Soheila Molaei, and David A Clifton. 2024. A brief review of hypernetworks in deep learning. Artificial Intelligence Review 57, 9 (2024), 1–29.
@@
- Ivison et al. (2022) Hamish Ivison, Akshita Bhagia, Yizhong Wang, Hannaneh Hajishirzi, and Matthew Peters. 2022. HINT: Hypernetwork Instruction Tuning for Efficient Zero-& Few-Shot Generalisation. arXiv preprint arXiv:2212.10315 (2022).
@@
- Litany et al. (2022) Or Litany, Haggai Maron, David Acuna, Jan Kautz, Gal Chechik, and Sanja Fidler. 2022. Federated learning with heterogeneous architectures using graph hypernetworks. arXiv preprint arXiv:2201.08459 (2022).